# A Priori Assessment of Rotational Invariance in Multiscale CNN-Based Subgrid-Scale Model for Wall-Bounded Turbulent Flows

### v1

@@ II.2 Dataset preparation
The dataset for the data-driven SGS model is obtained from the DNS data of the turbulent channel flow with the friction Reynolds number $\text{Re}_{\tau}=\frac{u_{\tau}\delta}{\nu}=180$. $u_{\tau}$ denotes the frictional mean velocity, $\nu$ is the kinematic viscosity, and $\delta$ is the channel half-width. The governing equations for incompressible continuity and Navier-Stokes equations are directly solved. A collocated grid with non-uniform spacing in the wall-normal ($y$) direction is used, whereby the no-slip boundary condition is applied to the walls and periodicity is applied in the streamwise ($x$) and spanwise ($z$) directions. The computational domains and computational grids are $(L_{x}\times L_{y}\times L_{z})=(4\pi\delta\times 2\delta\times 2\pi\delta)$ and $(N_{x}\times N_{y}\times N_{z})=(192\times 128\times 160)$, respectively. The unsteady solver employs a fractional-step method with second-order central differences for convective and viscous terms. The Adams-Bashforth method was adopted for time marching. The validation results (Fig. 2) show the wall-normal distributions of the mean streamwise velocity ($u^{+}$) and root-mean-square (rms) values of the fluctuating velocity components in the streamwise ($u_{\mathrm{rms}}^{+}$), wall-normal ($v_{\mathrm{rms}}^{+}$), and spanwise ($w_{\mathrm{rms}}^{+}$) directions, which are in good agreement with the results of Kim et al. kim1987 and wall law. Here, the superscript $(.)^{+}$ denotes scaling in wall units.
@@
The DNS flow fields are then box-filtered to separate the GS and SGS quantities, yielding the fDNS data. The GS variables are used as the inputs, whereas the SGS stress, $\tau_{ij}$, is calculated to provide the corresponding label data. As in the study by Jalaali &\& Okabayashijalaali2025, the datasets are comprised of 140 instantaneous fields, approximately 1,820 non-dimensional time units ($tU_{bulk}/\delta$), 25,180 viscous time scales ($tu_{\tau}^{2}/\nu$), and 291 flow throughs ($tU_{bulk}/L_{x}$), where $U_{bulk}$ denotes the bulk velocity. Here, the datasets are split 90:10 for training and testing.
@@ II.3 Baseline data-driven SGS model algorithm and framework
The DNN algorithm plays a crucial role owing to its ability to establish a functional relationship between the input features derived from the resolved flow fields and the six output variables $\tau_{ij}$. In this study, we adopted the multiscale-CNN model proposed by Jalaali and Okabayashi jalaali2025 with a minor change in the loss function (hereafter referred to as the MSC model) to represent multiple scales of a turbulent flow, which serves as the baseline architecture. We further developed two modified versions of this model, referred to as MSC2 and MSC3, in the subsequent sections to enhance rotational invariance. Jalaali &\& Okabayashijalaali2025 emphasized that the MSC model incorporated multiscale representation and physical processes across scales, resulting in more accurate predictions. Here, the MSC model is outlined briefly and depicted in Fig. 3. Based on the principle of energy cascade, where energy is transferred from larger to smaller scales, the input features were decomposed into separate scales representing large-, intermediate-, and small-scale eddies. The MSC model encodes the information progressively, starting from the largest scale and proceeding to the smallest. Here, we found that the use of three scales provided the most accurate predictions, offering an optimal tradeoff between model complexity and accuracy. In each stage, the encoded information from the previous scale is concatenated to form a comprehensive multiscale representation. This concatenation is important for integrating the inter-layer information without sacrificing detail, as noted by Huang et al.huang2018.
@@
where $w$ denotes the DNN weight or convolutional kernel in CNN; BN and $b$ denote batch normalization and bias, respectively. Here, $(.)_{l}$ denotes the hidden layer index (l=05)(l=0-5) and $(.)_{k}$ is the scale layer index; $k=1,2$, and $3$ represent the quarter-, half-, and full-scale layers, respectively. Therefore, the output $\tau^{p}_{ij}$ is $x_{3,5}$. 3D convolutional kernels with a uniform size of 3 in each spatial direction are employed in the CNN to represent the nature of turbulent flow interactions. The exponential linear unit (ELU) activation function $\varphi$ is applied, which introduces nonlinearity through a fixed mathematical mapping that transforms the datagoodfellow2016. ELU offers the advantage of a smooth gradient during backpropagation by allowing negative outputs, thereby helping to mitigate the vanishing gradient problemClevert2015Fast; kim2020. This property is particularly beneficial in capturing both the positive and negative components of the input variables. $\bar{D}_{ij}$ is chosen as the input variable, as it satisfies both Galilean and rotational invariancekajishima2017. Hence, $X_{1,0}$ is $\bar{D}_{ij}$. As emphasized by various researchers, such as SpalartSpalart2023, Brenner et al. brener2024, Wu et al. wu1028, and Prakash et al. prakash2022, the adherence of the input to the material objectivity requirement is critical for constructing data-driven SGS models
@@ II.4 Data-driven SGS model with rotational invariance
In the following, we describe the proposed model and its distinction from the baseline MSC model to satisfy the principle of material objectivity described in Eq. (1). We first note that relying solely on invariant input variables is not sufficient to ensure material objectivity, because the DNN operates as a composite function, as formulated in Eq. (6). As noted by Mo &\& ZhaoMo and Jaderberg et al.Jaderberg, CNN models have limitations in handling rotational orientation, requiring modification of the model. Here, in contrast to the MSC model, both bias and BN terms were excluded, as follows:
@@
This modification provides the data-driven SGS model with raw data representations, unaltered by normalization or bias, following the approach of Bin et al.Bin2022 and Cho et al. Cho2024 Physically, the inclusion of a bias term is unnecessary because the SGS stress output should vanish when the input tensor is zeroBin2022; Cho2024. In addition, these studies showed that removing the bias and BN improved the generalization of the data-driven SGS model, which enabled the model to make predictions beyond the training dataBin2022; Cho2024. Although BN is intended to improve the stability and convergence during training by independently normalizing the output of the CNN layers across mini-batchesioffe2015batch, we argue that this partitioning of the mini-batches breaks the spatial coherence of the data and disturbs the underlying physical representation of the vortex interaction. It is important to note that such a modification may result in an unphysical representation, as the local interactions are correlated in turbulent flows. As a result, the data-driven SGS model becomes more fitted to the training dataset and may lose its ability to generalize under rotational transformations owing to the disrupted spatial correlation across the mini-batches. We also provide a separate discussion on the effects of removing bias and BN in Appendix A. Therefore, by excluding the bias and BN terms while preserving the data structure through ELU activation, the convolutional kernel can be regarded as translationally invariant goodfellow2016. However, rotation introduces a different type of transformation that is not automatically considered by the model. Accordingly, we hypothesize that rotation transformation takes place only at the initial input layer. As the input is $\bar{D}_{ij}$ and the rotational transformation is applied prior to the first convolutional layer, the effect of rotation is confined to the input stage. Because $\bar{D}_{ij}$ is an objective quantity, such transformations do not compromise the extracted features. Moreover, the ELU activation preserves both positive and negative values of the input, ensuring that no data are discarded and the underlying physical behavior of the features is not distorted. This newly proposed model configuration is hereinafter referred to as MSC2.
@@
To further enhance the invariance properties and accuracy of the data-driven SGS model, we incorporated an STNJaderberg. Specifically, the MSC2 model was modified by adding the STN algorithm (Fig. 4), hereinafter referred to as the MSC3 model (Fig. 5). The main intention was to equip the data-driven SGS model with a transformation coordinate framework to help satisfy the principle of material objectivity. This approach is based on the assumption that, when given rotated inputs, a DNN algorithm that is sensitive to orientation may alter the spatial representation of the data regardless of the input quantity. The data-driven SGS model considerably lacks a spatial coordinate because neither the input nor the label data contain explicit coordinate information. Consequently, its ability to interpret the orientations of the flow structures is limited. Therefore, the STN algorithm is introduced to infer the geometric transformations and ensure spatial consistency. A concise description of the STN algorithm is provided here for completeness. A detailed explanation of its architecture and implementation can be found in the study by Jaderberg et al.Jaderberg
@@
The STN algorithm consists of three main components: a localization network, grid generator, and sampler. The localization network first predicts the transformation parameters $(.)_{\theta}$ that describe geometric operations such as rotation or translation from the rotated and/or translated input feature. It is implemented as a regular CNN whose weights are optimized via backpropagation using the loss propagated from the data-driven SGS model (Eq. (9)), as illustrated in Fig. 5. The loss function is minimized during training; a lower loss value indicates a more accurate prediction of the orientation, thereby allowing the network to infer the spatial transformation required to realign the input features. Through this process, the localization network learns to recognize the orientation of the input features. The predicted parameters $(.)_{\theta}$ are then passed to the grid generator, which provides a rectilinear grid of $G=(x_{i}^{t},y_{i}^{t},z_{i}^{t})$ corresponding to the $i$-th grid point that represents the grid coordinates in the target coordinate space of the STN output, where each grid point corresponds to a location to be sampled from the input features. Here, $G$ consists of normalized coordinate values. The coordinate transformation between the output grid $G$ and $(.)_{\theta}$ is performed using an affine transformation, denoted by $\mathcal{T}_{(.)}$, which allows rotation and translation. The affine transformation for the 3D domain is defined as followsJaderberg:
@@
Figure 4: Schematic of STN algorithm based on Ref.Jaderberg
By applying the affine transformation to $G$, the grid positions in the input feature are obtained as $(x_{i}^{s},y_{i}^{s},z_{i}^{s})$, which denote the source coordinates corresponding to the $i$-th grid point. These coordinates represent the coordinate information of the given input feature. Because the source coordinates generally do not align with the discrete grid points on the output feature grid, a sampler is employed to estimate the feature values at those locations. The sampler performs trilinear interpolation over the neighboring voxels surrounding each source coordinate to compute the interpolated value corresponding to the output feature grid. As illustrated in Fig. 6, the red arrows represent the flow of the sampler operation, whereas the black dashed line represents the mapping between the input and output features through the source coordinates. This ensures that the mapping between the input and output features remains spatially consistent Jaderberg. As a result, the STN algorithm transforms the rotated and/or translated input features into spatially aligned (non-rotated) representations, allowing the data-driven SGS model to interpret the flow structures consistently across different orientations. This capability is particularly critical in turbulence modeling, where the accurate recognition of the spatial orientation and coherence directly affects the representation of the vortex interactions. This configuration is expected to enhance the invariance property of the data-driven SGS model. The model configurations and comparisons are summarized in Table 1.
@@ II.5 Training of data-driven SGS model
The total loss function $\mathcal{L}$ consists of three components: the prediction loss ($\mathcal{L}_{d}$), mean loss ($\mathcal{L}_{m}$), and regularization loss ($\mathcal{L}_{r}$), as shown in Eq. (9). Here, the loss function proposed by Jalaali &\& Okabayashijalaali2025 for the MSC model is extended. An additional mean-loss term $\mathcal{L}_{m}$ is introduced to incorporate the wall-normal distribution information and improve the predictive accuracy in the near-wall region by following the approach of Arranz et al.arranz2024 Accordingly, $\gamma_{ij}$ is introduced to amplify the influence of each $\tau_{ij}$ component, thereby facilitating more effective updates to the kernel weightsjalaali2025. The values of $\gamma_{ij}$ are selected based on the global average magnitude of the respective $\tau_{ij}$ components, and further optimized through a hyperparameter tuning process. The selected values are $\gamma_{11}=1$, $\gamma_{22}=10$, $\gamma_{33}=10$, $\gamma_{12}=10$, $\gamma_{13}=10$, and $\gamma_{23}=100$. A regularization term $\mathcal{L}_{r}$ is included to prevent overfitting, with a regularization constant $\lambda$ set to $10^{-4}$. Here, $w_{k,l}$ denotes the convolutional kernel coefficients at scale $k$ and layer $l$.
To minimize the loss, we employ the adaptive moment estimation (ADAM) optimization algorithm kingma2017, which iteratively updates the convolutional kernel in the negative direction of the total loss gradient. Additionally, we apply a learning rate scheduler, with an initial learning rate of $10^{-3}$, which is reduced by a factor of 20 every 50 epochs. Here, an epoch refers to a complete pass through the entire training dataset. The data-driven SGS models were trained for 400 epochs with a mini-batch size of 8. The training was conducted using the PyTorch 1.13.1 machine learning library, and the computations were performed on a large-scale computer system consisting of eight NVIDIA A100 GPU cores, hosted at the D3 Center of The University of Osaka.
@@ III Verification of material objectivity
First, the evaluation focused on $\tau_{12}$, which is considered the most significant component in wall-bounded turbulent flows bose2024. The wall-normal distribution of the shear stress component $\tau_{12}$ is depicted in Fig. 7. Here, the non-rotated input implies $0^{\circ}$ rotation, and the same expression is used hereinafter. For the non-rotated input, the overall agreement between the MSC model and fDNS data is comparable to that between the MSC2 and MSC3 models, although the MSC3 model exhibits a better prediction capability in the near-wall regions. In the region of $\textit{y}^{+}>20$, the MSC model can reasonably capture the overall trend of the fDNS distribution, whereas the MSC2 and MSC3 models exhibit a slight underprediction and overprediction, respectively. As also discussed in Appendix A, this is likely attributed to the presence of bias and BN, allowing data-driven SGS model to fit the data distribution better. Despite the exclusion of both bias and BN, the results of the MSC2 and MSC3 models are comparable to those of fDNS. This indicates that removing bias and BN does not inherently degrade the model accuracy. For the rotated input, the MSC model shows significant deviations and is unable to predict accurately, which demonstrates that this model does not satisfy the requirement of material objectivity. In contrast, both the MSC2 and MSC3 models maintain consistent accuracy under rotation and show a slight deviation from the fDNS data. This result implies that the MSC2 and MSC3 models satisfy the requirement of objectivity or rotational invariance.
@@
Furthermore, we evaluated the prediction performance on $\tau_{11}$, which quantifies the unresolved stresses associated with the streamwise velocity fluctuations. $\tau_{11}$ represents the streamwise component of the SGS stress, reflecting the energy transfer between the resolved and subgrid-scale kinetic energy in the main flow directionsagaut2006. Following the same procedure as in the previous analysis, we assessed the performance of the data-driven SGS model for both non-rotated and rotated inputs. Fig. 9 shows the wall-normal distributions of $\tau_{11}$ for both non-rotated and rotated inputs. For the non-rotated input, all models successfully capture the distribution of $\tau_{11}$, showing excellent agreement with the fDNS data. In contrast, when the input fields are rotated, the MSC model fails to predict $\tau_{11}$ accurately. This discrepancy can be attributed to the sensitivity of $\tau_{11}$ to the input orientation. Unlike $\tau_{12}$, which represents a shear component between the upper and lower walls, $\tau_{11}$ might depend explicitly on the streamwise direction. Consequently, when the input is rotated, the streamwise direction changes its alignment relative to the learned coordinate system, making the prediction task significantly more difficult. Another possible factor is that $\tau_{11}$ exhibits a large value of the SGS stress, such that the rotation of the input field can lead to a more pronounced alteration in the corresponding feature representation. Nevertheless, both the MSC2 and MSC3 models are able to provide accurate predictions for the rotated input.
@@
Figure 10 presents the distribution of the number of occurrences $f$ for $\tau_{11}$ (top row) and $\tau_{12}$ (bottom row) at different rotation angles of $0^{\circ}$, $90^{\circ}$, and $270^{\circ}$. Here, the number of bins in the histogram is set to 200, which corresponds to the bin width of $0.01$. The $\tau_{11}$ and $\tau_{12}$ data within each bin are counted. $\tau_{11}$ shows a positive distribution with a pronounced right tail that indicates strong fluctuations. This positive contribution also significantly enhances the SGS turbulent kinetic energysagaut2006. The accurate prediction of $\tau_{11}$ can be crucial for maintaining the net forward energy transfer and SGS kinetic energy. The shear component $\tau_{12}$ takes both positive and negative values, reflecting the turbulent shear induced by the wall-normal velocity gradients, which is important in the momentum exchange of a turbulent flowdavidson2015. Underprediction of the value weakens the subgrid stress, whereas exaggeration of the tails results in overprediction of the shear.

### v2

@@ II.2 Dataset preparation
The dataset for the data-driven SGS model is obtained from the DNS data of the turbulent channel flow with the friction Reynolds number $\text{Re}_{\tau}=\frac{u_{\tau}\delta}{\nu}=180$. $u_{\tau}$ denotes the frictional mean velocity, $\nu$ is the kinematic viscosity, and $\delta$ is the channel half-width. The governing equations for incompressible continuity and Navier-Stokes equations are directly solved. A collocated grid with non-uniform spacing in the wall-normal ($y$) direction is used, whereby the no-slip boundary condition is applied to the walls and periodicity is applied in the streamwise ($x$) and spanwise ($z$) directions. The computational domains and computational grids are $(L_{x}\times L_{y}\times L_{z})=(4\pi\delta\times 2\delta\times 2\pi\delta)$ and $(N_{x}\times N_{y}\times N_{z})=(192\times 128\times 160)$, respectively. The unsteady solver employs a fractional-step method with second-order central differences for convective and viscous terms. The Adams-Bashforth method was adopted for time marching. The validation results (Fig. 2) show the wall-normal distributions of the mean streamwise velocity ($u^{+}$) and root-mean-square (rms) values of the fluctuating velocity components in the streamwise ($u_{\mathrm{rms}}^{+}$), wall-normal ($v_{\mathrm{rms}}^{+}$), and spanwise ($w_{\mathrm{rms}}^{+}$) directions, which are in good agreement with the results of Kim et al. [23] and wall law. Here, the superscript $(.)^{+}$ denotes scaling in wall units.
@@
The DNS flow fields are then box-filtered to separate the GS and SGS quantities, yielding the fDNS data. The GS variables are used as the inputs, whereas the SGS stress, $\tau_{ij}$, is calculated to provide the corresponding label data. As in the study by Jalaali &\& Okabayashi[21], the datasets are comprised of 140 instantaneous fields, approximately 1,820 non-dimensional time units ($tU_{bulk}/\delta$), 25,180 viscous time scales ($tu_{\tau}^{2}/\nu$), and 291 flow throughs ($tU_{bulk}/L_{x}$), where $U_{bulk}$ denotes the bulk velocity. Here, the datasets are split 90:10 for training and testing.
@@ II.3 Baseline data-driven SGS model algorithm and framework
The DNN algorithm plays a crucial role owing to its ability to establish a functional relationship between the input features derived from the resolved flow fields and the six output variables $\tau_{ij}$. In this study, we adopted the multiscale-CNN model proposed by Jalaali and Okabayashi [21] with a minor change in the loss function (hereafter referred to as the MSC model) to represent multiple scales of a turbulent flow, which serves as the baseline architecture. We further developed two modified versions of this model, referred to as MSC2 and MSC3, in the subsequent sections to enhance rotational invariance. Jalaali &\& Okabayashi[21] emphasized that the MSC model incorporated multiscale representation and physical processes across scales, resulting in more accurate predictions. Here, the MSC model is outlined briefly and depicted in Fig. 3. Based on the principle of energy cascade, where energy is transferred from larger to smaller scales, the input features were decomposed into separate scales representing large-, intermediate-, and small-scale eddies. The MSC model encodes the information progressively, starting from the largest scale and proceeding to the smallest. Here, we found that the use of three scales provided the most accurate predictions, offering an optimal tradeoff between model complexity and accuracy. In each stage, the encoded information from the previous scale is concatenated to form a comprehensive multiscale representation. This concatenation is important for integrating the inter-layer information without sacrificing detail, as noted by Huang et al.[18].
@@
where $w$ denotes the DNN weight or convolutional kernel in CNN; BN and $b$ denote batch normalization and bias, respectively. Here, $(.)_{l}$ denotes the hidden layer index (l=05)(l=0-5) and $(.)_{k}$ is the scale layer index; $k=1,2$, and $3$ represent the quarter-, half-, and full-scale layers, respectively. Therefore, the output $\tau^{p}_{ij}$ is $x_{3,5}$. 3D convolutional kernels with a uniform size of 3 in each spatial direction are employed in the CNN to represent the nature of turbulent flow interactions. The exponential linear unit (ELU) activation function $\varphi$ is applied, which introduces nonlinearity through a fixed mathematical mapping that transforms the data[16]. ELU offers the advantage of a smooth gradient during backpropagation by allowing negative outputs, thereby helping to mitigate the vanishing gradient problem[11, 24]. This property is particularly beneficial in capturing both the positive and negative components of the input variables. $\bar{D}_{ij}$ is chosen as the input variable, as it satisfies both Galilean and rotational invariance[22]. Hence, $X_{1,0}$ is $\bar{D}_{ij}$. As emphasized by various researchers, such as Spalart[42], Brenner et al. [7], Wu et al. [44], and Prakash et al. [37], the adherence of the input to the material objectivity requirement is critical for constructing data-driven SGS models
@@ II.4 Data-driven SGS model with rotational invariance
In the following, we describe the proposed model and its distinction from the baseline MSC model to satisfy the principle of material objectivity described in Eq. (1). We first note that relying solely on invariant input variables is not sufficient to ensure material objectivity, because the DNN operates as a composite function, as formulated in Eq. (6). As noted by Mo &\& Zhao[30] and Jaderberg et al.[20], CNN models have limitations in handling rotational orientation, requiring modification of the model. Here, in contrast to the MSC model, both bias and BN terms were excluded, as follows:
@@
This modification provides the data-driven SGS model with raw data representations, unaltered by normalization or bias, following the approach of Bin et al.[5] and Cho et al. [8] Physically, the inclusion of a bias term is unnecessary because the SGS stress output should vanish when the input tensor is zero[5, 8]. In addition, these studies showed that removing the bias and BN improved the generalization of the data-driven SGS model, which enabled the model to make predictions beyond the training data[5, 8]. Although BN is intended to improve the stability and convergence during training by independently normalizing the output of the CNN layers across mini-batches[19], we argue that this partitioning of the mini-batches breaks the spatial coherence of the data and disturbs the underlying physical representation of the vortex interaction. It is important to note that such a modification may result in an unphysical representation, as the local interactions are correlated in turbulent flows. As a result, the data-driven SGS model becomes more fitted to the training dataset and may lose its ability to generalize under rotational transformations owing to the disrupted spatial correlation across the mini-batches. We also provide a separate discussion on the effects of removing bias and BN in Appendix A. Therefore, by excluding the bias and BN terms while preserving the data structure through ELU activation, the convolutional kernel can be regarded as translationally invariant [16]. However, rotation introduces a different type of transformation that is not automatically considered by the model. Accordingly, we hypothesize that rotation transformation takes place only at the initial input layer. As the input is $\bar{D}_{ij}$ and the rotational transformation is applied prior to the first convolutional layer, the effect of rotation is confined to the input stage. Because $\bar{D}_{ij}$ is an objective quantity, such transformations do not compromise the extracted features. Moreover, the ELU activation preserves both positive and negative values of the input, ensuring that no data are discarded and the underlying physical behavior of the features is not distorted. This newly proposed model configuration is hereinafter referred to as MSC2.
@@
To further enhance the invariance properties and accuracy of the data-driven SGS model, we incorporated an STN[20]. Specifically, the MSC2 model was modified by adding the STN algorithm (Fig. 4), hereinafter referred to as the MSC3 model (Fig. 5). The main intention was to equip the data-driven SGS model with a transformation coordinate framework to help satisfy the principle of material objectivity. This approach is based on the assumption that, when given rotated inputs, a DNN algorithm that is sensitive to orientation may alter the spatial representation of the data regardless of the input quantity. The data-driven SGS model considerably lacks a spatial coordinate because neither the input nor the label data contain explicit coordinate information. Consequently, its ability to interpret the orientations of the flow structures is limited. Therefore, the STN algorithm is introduced to infer the geometric transformations and ensure spatial consistency. A concise description of the STN algorithm is provided here for completeness. A detailed explanation of its architecture and implementation can be found in the study by Jaderberg et al.[20]
@@
The STN algorithm consists of three main components: a localization network, grid generator, and sampler. The localization network first predicts the transformation parameters $(.)_{\theta}$ that describe geometric operations such as rotation or translation from the rotated and/or translated input feature. It is implemented as a regular CNN whose weights are optimized via backpropagation using the loss propagated from the data-driven SGS model (Eq. (9)), as illustrated in Fig. 5. The loss function is minimized during training; a lower loss value indicates a more accurate prediction of the orientation, thereby allowing the network to infer the spatial transformation required to realign the input features. Through this process, the localization network learns to recognize the orientation of the input features. The predicted parameters $(.)_{\theta}$ are then passed to the grid generator, which provides a rectilinear grid of $G=(x_{i}^{t},y_{i}^{t},z_{i}^{t})$ corresponding to the $i$-th grid point that represents the grid coordinates in the target coordinate space of the STN output, where each grid point corresponds to a location to be sampled from the input features. Here, $G$ consists of normalized coordinate values. The coordinate transformation between the output grid $G$ and $(.)_{\theta}$ is performed using an affine transformation, denoted by $\mathcal{T}_{(.)}$, which allows rotation and translation. The affine transformation for the 3D domain is defined as follows[20]:
@@
Figure 4: Schematic of STN algorithm based on Ref.[20]
By applying the affine transformation to $G$, the grid positions in the input feature are obtained as $(x_{i}^{s},y_{i}^{s},z_{i}^{s})$, which denote the source coordinates corresponding to the $i$-th grid point. These coordinates represent the coordinate information of the given input feature. Because the source coordinates generally do not align with the discrete grid points on the output feature grid, a sampler is employed to estimate the feature values at those locations. The sampler performs trilinear interpolation over the neighboring voxels surrounding each source coordinate to compute the interpolated value corresponding to the output feature grid. As illustrated in Fig. 6, the red arrows represent the flow of the sampler operation, whereas the black dashed line represents the mapping between the input and output features through the source coordinates. This ensures that the mapping between the input and output features remains spatially consistent [20]. As a result, the STN algorithm transforms the rotated and/or translated input features into spatially aligned (non-rotated) representations, allowing the data-driven SGS model to interpret the flow structures consistently across different orientations. This capability is particularly critical in turbulence modeling, where the accurate recognition of the spatial orientation and coherence directly affects the representation of the vortex interactions. This configuration is expected to enhance the invariance property of the data-driven SGS model. The model configurations and comparisons are summarized in Table 1.
@@ II.5 Training of data-driven SGS model
The total loss function $\mathcal{L}$ consists of three components: the prediction loss ($\mathcal{L}_{d}$), mean loss ($\mathcal{L}_{m}$), and regularization loss ($\mathcal{L}_{r}$), as shown in Eq. (9). Here, the loss function proposed by Jalaali &\& Okabayashi[21] for the MSC model is extended. An additional mean-loss term $\mathcal{L}_{m}$ is introduced to incorporate the wall-normal distribution information and improve the predictive accuracy in the near-wall region by following the approach of Arranz et al.[1] Accordingly, $\gamma_{ij}$ is introduced to amplify the influence of each $\tau_{ij}$ component, thereby facilitating more effective updates to the kernel weights[21]. The values of $\gamma_{ij}$ are selected based on the global average magnitude of the respective $\tau_{ij}$ components, and further optimized through a hyperparameter tuning process. The selected values are $\gamma_{11}=1$, $\gamma_{22}=10$, $\gamma_{33}=10$, $\gamma_{12}=10$, $\gamma_{13}=10$, and $\gamma_{23}=100$. A regularization term $\mathcal{L}_{r}$ is included to prevent overfitting, with a regularization constant $\lambda$ set to $10^{-4}$. Here, $w_{k,l}$ denotes the convolutional kernel coefficients at scale $k$ and layer $l$.
To minimize the loss, we employ the adaptive moment estimation (ADAM) optimization algorithm [25], which iteratively updates the convolutional kernel in the negative direction of the total loss gradient. Additionally, we apply a learning rate scheduler, with an initial learning rate of $10^{-3}$, which is reduced by a factor of 20 every 50 epochs. Here, an epoch refers to a complete pass through the entire training dataset. The data-driven SGS models were trained for 400 epochs with a mini-batch size of 8. The training was conducted using the PyTorch 1.13.1 machine learning library, and the computations were performed on a large-scale computer system consisting of eight NVIDIA A100 GPU cores, hosted at the D3 Center of the University of Osaka.
@@ III Verification of material objectivity
First, the evaluation focused on $\tau_{12}$, which is considered the most significant component in wall-bounded turbulent flows [6]. The wall-normal distribution of the shear stress component $\tau_{12}$ is depicted in Fig. 7. Here, the non-rotated input implies $0^{\circ}$ rotation, and the same expression is used hereinafter. For the non-rotated input, the overall agreement between the MSC model and fDNS data is comparable to that between the MSC2 and MSC3 models, although the MSC3 model exhibits a better prediction capability in the near-wall regions. In the region of $\textit{y}^{+}>20$, the MSC model can reasonably capture the overall trend of the fDNS distribution, whereas the MSC2 and MSC3 models exhibit a slight underprediction and overprediction, respectively. As also discussed in Appendix A, this is likely attributed to the presence of bias and BN, allowing data-driven SGS model to fit the data distribution better. Despite the exclusion of both bias and BN, the results of the MSC2 and MSC3 models are comparable to those of fDNS. This indicates that removing bias and BN does not inherently degrade the model accuracy. For the rotated input, the MSC model shows significant deviations and is unable to predict accurately, which demonstrates that this model does not satisfy the requirement of material objectivity. In contrast, both the MSC2 and MSC3 models maintain consistent accuracy under rotation and show a slight deviation from the fDNS data. This result implies that the MSC2 and MSC3 models satisfy the requirement of objectivity or rotational invariance.
@@
Furthermore, we evaluated the prediction performance on $\tau_{11}$, which quantifies the unresolved stresses associated with the streamwise velocity fluctuations. $\tau_{11}$ represents the streamwise component of the SGS stress, reflecting the energy transfer between the resolved and subgrid-scale kinetic energy in the main flow direction[38]. Following the same procedure as in the previous analysis, we assessed the performance of the data-driven SGS model for both non-rotated and rotated inputs. Fig. 9 shows the wall-normal distributions of $\tau_{11}$ for both non-rotated and rotated inputs. For the non-rotated input, all models successfully capture the distribution of $\tau_{11}$, showing excellent agreement with the fDNS data. In contrast, when the input fields are rotated, the MSC model fails to predict $\tau_{11}$ accurately. This discrepancy can be attributed to the sensitivity of $\tau_{11}$ to the input orientation. Unlike $\tau_{12}$, which represents a shear component between the upper and lower walls, $\tau_{11}$ might depend explicitly on the streamwise direction. Consequently, when the input is rotated, the streamwise direction changes its alignment relative to the learned coordinate system, making the prediction task significantly more difficult. Another possible factor is that $\tau_{11}$ exhibits a large value of the SGS stress, such that the rotation of the input field can lead to a more pronounced alteration in the corresponding feature representation. Nevertheless, both the MSC2 and MSC3 models are able to provide accurate predictions for the rotated input.
@@
Figure 10 presents the distribution of the number of occurrences $f$ for $\tau_{11}$ (top row) and $\tau_{12}$ (bottom row) at different rotation angles of $0^{\circ}$, $90^{\circ}$, and $270^{\circ}$. Here, the number of bins in the histogram is set to 200, which corresponds to the bin width of $0.01$. The $\tau_{11}$ and $\tau_{12}$ data within each bin are counted. $\tau_{11}$ shows a positive distribution with a pronounced right tail that indicates strong fluctuations. This positive contribution also significantly enhances the SGS turbulent kinetic energy[38]. The accurate prediction of $\tau_{11}$ can be crucial for maintaining the net forward energy transfer and SGS kinetic energy. The shear component $\tau_{12}$ takes both positive and negative values, reflecting the turbulent shear induced by the wall-normal velocity gradients, which is important in the momentum exchange of a turbulent flow[12]. Underprediction of the value weakens the subgrid stress, whereas exaggeration of the tails results in overprediction of the shear.